Three times an even interger is two less than 4 times the next lower even integer. Find the interger ?
Found 2 solutions by checkley71, Dalmo:
Answer by checkley71(8403)

3x=(x-2)-2
3x=x-4
3x-x=-4
2x=-4
x=-4/2
x=-2 answer.
proof:
3*-2=(-2-2)-2
-6=-6

Answer by Dalmo(2)

Word problems can be very confusing at times, so pay close attention. Obviously, the person who answered before me made a mistake. It's very important to read every word.
-
Three times an even interger is two less than 4 times the next lower even integer. Find the interger ?
-
Let's call the even integer x.
So we have three times that, or 3x
-
That is equal to:
Two less than ..., or ...-2
Four times..., so far 4(...)-2
The next lower even integer, so finally 4(x-2)-2
-
So this is the equation we get.
3x = 4(x-2)-2
3x = 4x -8 -2 (Distribute the 4)
3x = 4x -10 (Simplify)
-1x= -10 (Subtract 4x from both sides)
x = 10 (Divide by -1)
-
Proof:
3(10)= 4(10-2) -2
30 = 4(8) -2
30 = 32 -2
30 = 30 (True)
-
I hope that helps!
-
For more help, go to www.ziizoo.com/ref/dalmo
-
Get 30 minutes with any professional tutor for FREE.
-
I'll see you there
-
Dalmo